OpenStudy (anonymous):

Definite Integral on interval [0,1] of function (e^-x+1)/e^-x

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\frac{ e ^{-x}+1 }{ e ^{-x} }\]

jimthompson5910 (jim_thompson5910):

break up the fraction (A+B)/C = A/C + B/C

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

do you see how it simplifies?

OpenStudy (anonymous):

1+e^x

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

that's nice

OpenStudy (anonymous):

very nice

OpenStudy (anonymous):

e

OpenStudy (anonymous):

is the answer

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

thank you again!